A 98.0 N grocery cart is pushed 12.0 m along an aisle by a shopper who exerts a constant horizontal foce of 40.0 N. If all frictional forces are neglected and the cart starts from rest, what is the grocery cart's final speed?
1
Expert's answer
2012-02-09T08:20:53-0500
As the acceleration of gravity g = 9.8 m/s^2, then the mass of cart is
m =
98/9,8 = 10 kg.
Using the Newton's Second Law
F = ma. So, the cart's
acceleration is a = F/m = 40/10 = 4 m/s^2.
As the initial speed was zero, we
get the equation for displacement and final speed:
S = a*t^2/2, V = at.
t
= sqrt(2*S/a).
V = a*sqrt(2*S/a) = sqrt(2*S*a) = sqrt(2*12*4) = sqrt(96) =
9.8 m/s.